# DeepASA: An Object-Oriented Multi-Purpose Network for Auditory Scene Analysis

### v2

@@ 6 Conclusion
Limitations and future work The main limitation is the large parameter size of the ATST adopted for the SED decoder. In this work, the separated object features are reprocessed by the patch-based ATST to keep the advantage of ATST pre-trained on a large set of audio classification datasets. Future research could explore pre-training the novel classification layer designed to be compatible with object features, using a large audio dataset like ATST.
@@
This work was supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945) and STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT), the BK21 FOUR program through the NRF grant funded by the Ministry of Education of Korea government (MOE).
@@ Appendix A Detailed specifications of the SED decoder
In this section, we provide detailed specifications of the SED decoder. The SED decoder consists of three branches: ATST, T-CRNN, and F-CRNN. The specifications of the SED decoder are presented in Table 5. First, in the ATST branch, the object features are converted into patches of length $T_{p}=4,F_{p}=64$. The sequence length corresponds to the number of these transformed patches, and the patches are reshaped along the channel dimension. Since the pre-trained ATST takes a single-channel mel spectrogram as input, a linear layer is added before ATST to match the embedding dimension, followed by adaptive pooling to align the time frame to $T^{\prime}=40$. In the T-CRNN branch, seven convolution layers are used, as in [45], followed by a linear layer after concatenating the output of the ATST branch. The channel dimension and pooled frequency dimension are then reshaped into a single dimension, with the time dimension treated as the sequence dimension, and passed through a T-GRU. Finally, in the F-CRNN branch, pooling is focused on the time dimension, followed by two convolution layers, after which the channel dimension and pooled time dimension are reshaped into a single dimension, and the frequency dimension is treated as the sequence dimension before passing through an F-GRU. The combined output of the ATST and T-CRNN branches then passes through a fully connected layer with an output dimension of 1 to produce the timestamp output, a fully connected layer with an output dimension of $C=13$ (the number of classes) to produce the sound event map, and is subsequently combined with the output of the F-CRNN after passing through attentive statistics pooling (ASP) [40] and a fully connected layer, producing the final class output.
@@
patchification
$\frac{TF}{T_{p}F_{p}}\times DT_{p}F_{p}$
(3,3)@64, LN, Mish (3,3)@64, LN, Mish
@@ Appendix C Effect of direct and reverb decoder on DoAE
Furthermore, to evaluate the performance of dereverberation, we compared the results of reverberant, direct (dereverberation, $\text{SI\text{-}SDR}_{\mathbf{s}}$) and reverb source separation ($\text{SI\text{-}SDR}_{\mathbf{h}}$), as well as the speech-to-reverberation modulation ratio (SRMR) [55]. SRMR is only calculated for classes corresponding to male and female speech. The experimental results demonstrate that the proposed model performs well in dereverberation, achieving performance comparable to full source separation. Additionally, with an SRMR of 6.5, the model indicates strong dereverberation effectiveness.
@@ Appendix D Detailed analysis on time-varying learnable window
In this section, we present experiments to investigate the role and functionality of the time-varying learnable window. In the audio encoder, the short-time Fourier transform (STFT) is widely used to convert a waveform into a complex spectrogram. Unlike speech separation, which considers only the speech class, USS involves various classes, necessitating dynamic adjustments to the configuration of the window. Although many conventional methods rely on fixed windows [28, 12, 38], this reliance limits their performance [56]. To address this limitation, we propose a time-varying learnable window that can dynamically adjust the window configuration for both USS and ASA.
@@
We conducted a detailed ablation study on the effect of the time-varying learnable window. The results are presented in Table 7. Initially, we compared encoding methods using complex STFT and learnable 1D convolution kernels. The performance of complex STFT generally exceeded that of the 1D convolution kernel, suggesting that while the 1D convolution kernel extracts learnable features, complex STFT is more effective at explicitly capturing frequency information in the feature representation. Additionally, the time-invariant learnable window learns a fixed window across time, resulting in a window length that is similar to that of the STFT. The time-invariant learnable window can be obtained by performing mean pooling in the temporal dimension before being passed through the linear layers of the time-varying learnable window. Next, we examined TDY-CNN [56], which learns attention weights for each time frame to control how much information is passed through, capturing phoneme variation. However, this approach has the limitation of being unable to adjust the frequency band information. In contrast, the proposed method allows for control over the frequency band by adjusting the width of the main lobe.
@@
Figure 11 shows a comparison of the sample between conventional STFT and the proposed time-varying learnable window. Figure 11(a) shows the time-domain waveform of the audio mixture, and Figure 11(b) plots its frame-wise standard deviation over time. Figure 11(c), (d), and (e) present the SED result of conventional STFT, that of time-varying learnable window, and ground truth, respectively, with binary values between 0 and 1 (black indicates 0, white indicates 1). The results show that with conventional STFT, object 1 was not captured, and object 3 failed in both classification and timestamp estimation. However, using the proposed method, the model can detect object 1, which improves the sound event detection performance of object 3.
Finally, Figure 12 shows a GradCAM [57] analysis of the separated object features immediately after the object splitter, to examine how well the object features encapsulate the properties of the corresponding sound objects. The model with conventional STFT struggled with object separation, as regions of speech are still embedded within telephone regions. In contrast, the proposed method more effectively isolates the telephone regions by suppressing most of the speech-related components. This demonstrates that the proposed approach effectively separates sound objects based on auditory information.
@@ Appendix F Comparison with SOTA models for ASA2 dataset and real-world demonstration
Comparison USS and SELD performance on ASA2 dataset We compared the performance of the proposed DeepASA model on the ASA2 dataset with that of other SOTA models in Table 8. First, the proposed model achieved superior USS performance compared to other models [26, 46, 48, 38], even though it operates with a MIMO setup and performs a dereverberation task. Despite having a lightweight feature aggregation module compared to DeFT-Mamba, the model still achieved SOTA performance. This improvement is likely due to the ability of the model to utilize various auditory information, such as class, timestamp, DoA, and noise, enabling better separation of foreground sources. Next, in terms of SELD performance, the proposed model significantly outperformed existing models. This is because ASA2 consists entirely of polyphonic audio, and earlier approaches that estimate the outputs directly from mixtures are unsuitable for such overlapped sources. Our model avoids this limitation by separating the objects.
@@
ByteDance-uss [7] 5.2 8.3 - - - - - 28 (+80.7) M 40.1 G
TF-GridNet [46] 8.7 10.7 - - - - - 14.7 M 462 G
SpatialNet [48] 9.6 10.2 - - - - - 7.3 M 71.8 G
DeFT-Mamba [38] 10.4 11.3 - - - - - 3.6 M 83.8 G
EINV2 [52] - - 48.5 39.5 27.1 51.5 0.431 51.5 M 6.7 G
ResNet Conformer [58] - - 45.7 41.0 26.6 53.7 0.414 13.6 M 7.6 G
SELD-Mamba [54] - - 43.5 42.7 25.5 56.7 0.396 75.1 M 4.3 G
MFF-EINV2 [49] - - 42.1 43.2 25.8 60.7 0.381 54.8 M 14.3 G

### v3

@@ 6 Conclusion
Limitations and future work The main limitation is the large parameter size of the ATST adopted for the SED decoder. In this work, the separated object features are reprocessed by the patch-based ATST to keep the advantage of ATST pre-trained on a large set of audio classification datasets. For the aspect of the bias of the dataset, the reverberation time was set between 0.2 and 0.6 seconds, so the performance of the model may degrade in environments with reverberation times longer than 0.6 seconds. Additionally, because the background noise was simulated with an SNR range of 6 to 30 dB, the performance of the model is also expected to decline when the background noise has an SNR below 0 dB and is louder than the foreground source. Next, regarding the potential misuse, since this model can extract speech of individual speakers and analyze their types and directions, there is a concern that it could be exploited for eavesdropping.
Future research could explore pre-training the novel classification layer designed to be compatible with object features, using a large audio dataset like ATST.
@@
This work was supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945) and STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT), the BK21 FOUR program through the NRF grant funded by the Ministry of Education of Korea government (MOE), and the Center for Applied Research in Artificial Intelligence (CARAI) funded by DAPA and ADD (UD230017TD).
@@
  • Appendix G presents the experimental results along with the number of sound sources.
@@ Appendix A Detailed specifications of the SED decoder
In this section, we provide detailed specifications of the SED decoder. The SED decoder consists of three branches: ATST, T-CRNN, and F-CRNN. The specifications of the SED decoder are presented in Table 5. First, in the ATST branch, the object features are converted into patches of length $T_{p}=4,F_{p}=64$. The sequence length corresponds to the number of these transformed patches, and the patches are reshaped along the channel dimension. Since the pre-trained ATST takes a single-channel mel spectrogram as input, a linear layer is added before ATST to match the embedding dimension, followed by adaptive pooling to align the time frame to $T^{\prime}=40$. In the T-CRNN branch, seven convolution layers are used, as in [schmid2024improving], followed by a linear layer after concatenating the output of the ATST branch. The channel dimension and pooled frequency dimension are then reshaped into a single dimension, with the time dimension treated as the sequence dimension, and passed through a T-GRU. Finally, in the F-CRNN branch, pooling is focused on the time dimension, followed by two convolution layers, after which the channel dimension and pooled time dimension are reshaped into a single dimension, and the frequency dimension is treated as the sequence dimension before passing through an F-GRU. The combined output of the ATST and T-CRNN branches then passes through a fully connected layer with an output dimension of 1 to produce the object activation curve, a fully connected layer with an output dimension of $C=13$ (the number of classes) to produce the sound event map, and is subsequently combined with the output of the F-CRNN after passing through attentive statistics pooling (ASP) [okabe2018attentive] and a fully connected layer, producing the final class output.
@@
patchification $\frac{TF}{T_{p}F_{p}}\times DT_{p}F_{p}$ (3,3)@64, LN, Mish (3,3)@64, LN, Mish
@@ Appendix C Effect of direct and reverb decoder on DoAE
Furthermore, to evaluate the performance of dereverberation, we compared the results of reverberant, direct (dereverberation, $\text{SI\text{-}SDR}_{\mathbf{s}}$) and reverb source separation ($\text{SI\text{-}SDR}_{\mathbf{h}}$), as well as the speech-to-reverberation modulation ratio (SRMR) [falk2010non]. SRMR is only calculated for classes corresponding to male and female speech. The experimental results demonstrate that the proposed model performs well in dereverberation, achieving performance comparable to full source separation. Additionally, with an SRMR of 6.5, the model indicates strong dereverberation effectiveness.
@@ Appendix D Detailed analysis on time-varying learnable window
In this section, we present experiments to investigate the role and functionality of the time-varying learnable window. In the audio encoder, the short-time Fourier transform (STFT) is widely used to convert a waveform into a complex spectrogram. Unlike speech separation, which considers only the speech class, USS involves various classes, necessitating dynamic adjustments to the configuration of the window. Although many conventional methods rely on fixed windows [adavanne2018sound, shao2024fine, lee2025deft], this reliance limits their performance [kim2022temporal]. To address this limitation, we propose a time-varying learnable window that can dynamically adjust the window configuration for both USS and ASA.
@@
We conducted a detailed ablation study on the effect of the time-varying learnable window. The results are presented in Table 7. Initially, we compared encoding methods using complex STFT and learnable 1D convolution kernels. The performance of complex STFT generally exceeded that of the 1D convolution kernel, suggesting that while the 1D convolution kernel extracts learnable features, complex STFT is more effective at explicitly capturing frequency information in the feature representation. Additionally, the time-invariant learnable window learns a fixed window across time, resulting in a window length that is similar to that of the STFT. The time-invariant learnable window can be obtained by performing mean pooling in the temporal dimension before being passed through the linear layers of the time-varying learnable window. Next, we examined TDY-CNN [kim2022temporal], which learns attention weights for each time frame to control how much information is passed through, capturing phoneme variation. However, this approach has the limitation of being unable to adjust the frequency band information. In contrast, the proposed method allows for control over the frequency band by adjusting the width of the main lobe.
@@
Figure 11 shows a comparison of the sample between conventional STFT and the proposed time-varying learnable window. Figure 11(a) shows the time-domain waveform of the audio mixture, and Figure 11(b) plots its frame-wise standard deviation over time. Figure 11(c), (d), and (e) present the SED result of conventional STFT, that of time-varying learnable window, and ground truth, respectively, with binary values between 0 and 1 (black indicates 0, white indicates 1). The results show that with conventional STFT, object 1 was not captured, and object 3 failed in both classification and activation estimation. However, using the proposed method, the model can detect object 1, which improves the sound event detection performance of object 3.
Finally, Figure 12 shows a GradCAM [selvaraju2017grad] analysis of the separated object features immediately after the object splitter, to examine how well the object features encapsulate the properties of the corresponding sound objects. The model with conventional STFT struggled with object separation, as regions of speech are still embedded within telephone regions. In contrast, the proposed method more effectively isolates the telephone regions by suppressing most of the speech-related components. This demonstrates that the proposed approach effectively separates sound objects based on auditory information.
@@ Appendix E Detailed study on Chain-of-Inference
Table 8: Performance comparison by repeating chain-of-inference
Stage USS SED DoAE SELD $\downarrow$ Complexities
without CoI 11.0 11.7 28.8 70.2 18.5 76.9 0.230 8.2 (+96.8) M 99.1 G
[…]
CoI (2nd stage) 11.1 11.9 26.8 72.3 17.2 77.5 0.216 16.0 (+96.8 M) 118.9 G
Additionally, the CoI mechanism utilizes the high-level information (DoA, SED) to refine the object feature separation based on the temporal coherence matching (TCM), so using the same mechanism multiple times can bias the results towards the direction of prioritizing TCM. We present the results of performing the chain-of-inference multiple times as Table 8. As a result, using CoI only once resulted in the highest performance.
@@ Appendix F Comparison with SOTA models for ASA2 dataset and real-world demonstration
Comparison USS and SELD performance on ASA2 dataset We compared the performance of the proposed DeepASA model on the ASA2 dataset with that of other SOTA models in Table 9. First, the proposed model achieved superior USS performance compared to other models [kong2023universal, wang2023tf, quan2024spatialnet, lee2025deft], even though it operates with a MIMO setup and performs a dereverberation task. Despite having a lightweight feature aggregation module compared to DeFT-Mamba, the model still achieved SOTA performance. This improvement is likely due to the ability of the model to utilize various auditory information, such as class, activation, DoA, and noise, enabling better separation of foreground sources. Next, in terms of SELD performance, the proposed model significantly outperformed existing models. This is because ASA2 consists entirely of polyphonic audio, and earlier approaches that estimate the outputs directly from mixtures are unsuitable for such overlapped sources. Our model avoids this limitation by separating the objects.
@@
ByteDance-uss [kong2017joint] 5.2 8.3 - - - - - 28 (+80.7) M 40.1 G
TF-GridNet [wang2023tf] 8.7 10.7 - - - - - 14.7 M 462 G
SpatialNet [quan2024spatialnet] 9.6 10.2 - - - - - 7.3 M 71.8 G
DeFT-Mamba [lee2025deft] 10.4 11.3 - - - - - 3.6 M 83.8 G
EINV2 [cao2021improved] - - 48.5 39.5 27.1 51.5 0.431 51.5 M 6.7 G
ResNet Conformer [niu2023experimental] - - 45.7 41.0 26.6 53.7 0.414 13.6 M 7.6 G
SELD-Mamba [mu2024seld] - - 43.5 42.7 25.5 56.7 0.396 75.1 M 4.3 G
MFF-EINV2 [mu2024mff] - - 42.1 43.2 25.8 60.7 0.381 54.8 M 14.3 G
[…]
@@
Additionally, we conducted an experiment by fine-tuning the pretrained DeepASA over datasets generated from real-world room impulse responses (RIRs). The fine-tuned model was then evaluated by unseen real-world RIRs. From Table 10, the experimental results show a slight decrease in USS performance, as well as in most of SELD performance. This performance degradation can be attributed to the discrepancy in the hidden representation of simulation and real RIRs. Ideally, training with a large set of RIRs, including both simulated and measured RIRs, would improve the generalizability of models.
Table 10: Comparison between datasets with simulated and real-world RIRs
RIR USS SED DoAE SELD $\downarrow$
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$
simulated RIR 11.0 11.7 28.8 70.2 18.5 76.9 0.230
real RIR 10.6 11.4 32.7 66.4 13.4 72.3 0.254
Appendix G Experimental results along with the number of sound sources
We performed inference when the number of sound sources was within the maximum estimable range [2, 5]. and when it exceeded the maximum estimable range [6, 7]. The experimental results are presented in Table 11. When the number of foreground sources is 6 or 7, resulting in a sharp decline in performance. The experimental results show that when there are 6 sources, two sources are not separated, leading to very low performance only for the two sources. When there are 7 sources, three sources are mixed and mapped to one track in most cases, resulting in very low performance for three of the sources.
Table 11: Experimental results along with the number of sound sources
Number of sources USS SED DoAE SELD $\downarrow$
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$
2 17.7 17.1 22.1 82.4 10.5 88.5 0.143
3 14.6 14.5 22.4 80.8 13.1 86.3 0.160
4 8.5 10.6 33.8 65.6 20.9 72.5 0.269
5 7.9 10.3 34.9 64.5 22.1 71.3 0.280
6 Total 2.4 5.9 47.7 48.5 40.3 54.5 0.418
Top 4 average 7.0 9.6 38.2 61.2 25.6 66.7 0.311
Bottom 2 average -6.8 -1.5 69.1 23.1 69.7 30.1 0.636
7 Total 0.2 4.5 59.0 36.6 51.7 42.4 0.521
Top 4 average 6.6 9.1 38.9 60.6 29.2 64.5 0.325
Bottom 2 average -8.3 -1.6 85.8 4.6 81.7 12.9 0.784